OpenStudy (anonymous):

A boy whirls a stone with mass 250g in a horizontal circle of radius 1.5m and at height 3.0 m above level ground. The string breaks, and the stone flies off horizontally and strikes the ground after travelling a horizontal distance of 8m. What is the magnitude of the centripetal acceleration and the centripetal force of the stone during circular motion?

OpenStudy (anonymous):

D= (1/2)at^2+(v(initial)*t) 3 = .5*9.8t^2 + 0 t= .2474 s horizontal velocity = 8m / .2474 s = 32.3363 m/s cent. acc. = v^2/r cent. acc. = 32.3363/1.5 = 697.091 m/s^2 <----Answer F = ma F=.250*697.091= F=174.273 kg * m/s^2 or N <----Answer